# Song Form-aware Full-Song Text-to-Lyrics Generation with Multi-Level Granularity Syllable Count Control

### v1

@@ Abstract
Song Form-aware Full-Song Text-to-Lyrics Generation with
Multi-Level Granularity Syllable Count Control
Yunkee Chae1, Eunsik Shin2, Hwang Suntae2, Seungryeol Paik2, Kyogu Lee1,2,3 1IPAI, 2Department of Intelligence and Information, 3AIIS, Seoul National University {yunkimo95,esshin,iamsuntae1,paik402,kglee}@snu.ac.kr
1 Introduction
The field of lyrics information processing Watanabe and Goto (2020) presents unique challenges that extend beyond traditional text generation, as lyrics must align with both song form structures (such as verses, choruses, and bridges) and specific syllabic constraints. While natural language generation models have shown promise in generating coherent text, applying these models to lyrics introduces additional complexities due to the need for musical and rhythmic alignment.
Several studies have made notable progress in aligning lyrics with melodies through melody-aligned or midi-aligned generation methods Watanabe et al. (2018); Lu et al. (2019); Sheng et al. (2021). These efforts have been valuable in addressing melody-lyrics synchronization. However, the scarcity of large-scale audio-lyrics aligned datasets still poses a challenge for these approaches, limiting their generalizability across diverse musical forms Meseguer-Brocal et al. (2019); Yu et al. (2021); Durand et al. (2023). Given this data limitation, we concentrate on a crucial aspect of melody-lyrics alignment: the syllable count, which directly impacts the rhythmic fit of lyrics within a musical composition. However, precise syllable control is challenging, as models like GPT Brown et al. (2020), which rely on sub-word tokens, may struggle to manage syllables effectively.
@@ 1 Introduction
2 Multi-Level Granularity Lyrics Generation
Our system supports flexible generation across different levels of textual granularity (words, phrases. lines, and paragraphs), ensuring precise syllable control at each level.
The model is conditioned on input text, song structure (e.g., verses, choruses), and syllable constraints, generating lyrics that adhere to these requirements.
Token Structure.
To facilitate the generation of lyrics aligned with specific song forms, we define a structured token system.
Song form tokens (e.g., <VERSE>, <CHORUS>) are used to signify transitions between different parts of the song.
Following the song form token, we introduce a syllable count token, <SYL:>, where denotes the designated syllable count for the section, inspired by Kim et al. (2023a) and Guo et al. (2022).
These tokens help the model generate text at varying levels of granularity while maintaining syllabic precision.
@@ 2 Multi-Level Granularity Lyrics Generation
Inspired by the ideas presented in Bai et al. (2021), we implement distinct end tokens for each level of granularity in our model. Our model utilizes a structured approach with specific tokens to generate lyrics at different levels of granularity. For generating a complete paragraph, the <GEN_P> followed by the syllable count token and paragraph, and concluded with an <END_P>. To simulate more detailed generation control, the granularity choice extends to generating entire lines or lines segmented into phrases and words. Deciding to generate an entire line involves adding a <GEN_L> token at the beginning and concluding with an <END_L> token, similar to paragraph generation.
Alternatively, suppose a line is to be created with phrases and words. In that case, <GEN_L_NW> tokens are used along with respective syllable count tokens to indicate the start of such detailed line generation.
Each segment, whether phrases or words, begins with <GEN_N> or <GEN_W> tokens, respectively, followed by the corresponding syllable count token and text, and concluded with <END_NW> tokens. Upon completing a line detailed with phrases and words, an <END_L> token signifies its end.
An illustration of the generation plan is provided in Figure 2.
Semantic Embedding for Input Text.
We employ the SentenceTransformer111https://huggingface.co/sentence-transformers/all-mpnet-base-v2 Reimers and Gurevych (2019) model to capture the semantic content of lyrics, a method inspired by the approach of Watanabe et al. Watanabe and Goto (2023), who used Stable Diffusion Rombach et al. (2022) and pre-trained Vision Transformer Dosovitskiy et al. (2020) for semantic extraction in lyrics.
We use semantic embeddings of full lyrics as a substitute condition for input text during training, enabling the model to generate relevant lyrics by analyzing the semantic condition of any input text provided during inference.
Inference.
During the inference, we initiate the process with the provided semantic content embedding from the input text and proceed to decode each token in an autoregressive manner.
At each step of decoding we provide the model with special tokens that align with our predefined generation plan, rather than having the model predict these special tokens – such as song form, syllable counts, and generation directives tokens (<GEN_*>).
We continue to generate lyrics corresponding to a condition until an end for the specified granularity (<END_*>) is predicted.
Following this, the next set of condition tokens is given to the model.
This procedure repeats, ensuring every condition token in the generation plan is systematically applied, guiding the generation process to completion.
3 Results
In this section, we present our experimental results, including preliminary findings from recent large LLM models, results from our generation model, and the song form consistency of lyrics generated by our model. Details of experimental setups and evaluation metrics are provided in Appendix C.
3.1 Preliminary Experiment
Large language models (LLMs) like ChatGPT Ouyang et al. (2022) have shown remarkable performance in natural language generation recently. However, these models often struggle with accurately counting syllables Sun et al. (2023). To evaluate the performance of LLMs in generating content conditioned on syllable count and structure, we conducted a simple experiment. In this test, we task an LLM to create a verse by providing input text along with the number of lines and the syllable count for each line. A trial is considered a failure if the model does not produce the correct number of paragraphs or lines within five attempts. The success rates are approximately 38% for ChatGPT 3.5 and 57% for ChatGPT 4.0. Table 1 shows the results for the 688 successful samples from both ChatGPT 3.5 and ChatGPT 4.0. The results from Table 1 indicate that while LLMs are capable of generating lyrics that are contextually appropriate to the input text, they perform poorly in matching the required syllable count.
3.2 Lyrics Generation
In Table LABEL:tab:result_gen, we present the performance of our models across various settings. For each training method, we report on models trained from scratch as well as those initialized with large-scale pre-trained GPT-2 weights.
Front models serve as the baselines, where all generation plans are established at the start, before the <LYR_START> token, allowing for the production of complete lyrics without the guidance of generation plans during the generation process. In other words, we concatenate the special tokens and use them as a prompt to generate the output lyrics. Back introduces the generation plan directly to the model during the actual generation process, as illustrated in Figure 2. Both merges the strategies of Front and Back. It begins by listing all generation plans before <LYR_START>, then proceeds to generate lyrics using the Back approach. -S and -P at the end of the name of each model refer to whether the models are trained from scratch or from a pre-trained model, respectively.
All models carry some risk of failure during the generation process. For instance, if a model continues generating text without producing the <END_*> token, it will run until reaching the maximum length limit, resulting in a failed attempt. Additionally, in the Front method, there is a risk that the generated output may not align with the planned structure. The failure rates for each model are detailed in the Appendix D.
Table LABEL:tab:result_gen presents the metric values calculated from the lyrics successfully generated by all models. Only samples that are successfully generated by all models within 10 attempts, for each input condition in the evaluation set, are included. One-tailed Wilcoxon signed-rank tests are performed on a per-generation plan basis to test directional hypotheses. The BERT-Score (BERT-S) Zhang et al. (2019) is calculated between the input text and the generated lyrics. The average BERT-S between the input text and the original lyrics in the test set is 0.799, which serves as an upper bound.
Table LABEL:tab:result_gen shows that pre-trained models outperform those trained from scratch in terms of perplexity (PPL), with the Front and Both models showing the most favorable PPL results. A high PPL indicates that the model is experiencing significant confusion. The reason Front and Both models exhibit lower PPL compared to Back models is likely because the model can preview all conditions before generating lyrics, enabling it to make more informed predictions. In contrast, the Back models, which apply conditioning at each decoding step, have limited foresight regarding upcoming conditions, potentially leading to higher PPL. This limitation also implies that the Back models cannot fully anticipate the entire structure of the lyrics including the end of lyrics. However, given conditions directly at each decoding step could confer advantages in predicting syllable counts. The Both models attempt to integrate the strengths of both Front and Back models but at the expense of increased sample length. Thus, each conditioning approach inherently involves a trade-off.
On the other hand, for syllable count distance (SCD) and syllable count error rate (SCErr), models trained from scratch consistently excel over their pre-trained counterparts. A possible reason for these results is that the scratch models appear to learn more flexibly regarding special tokens, such as syllable tokens, compared to the pre-trained models. For all models, the metrics related to syllable counts perform poorer at the paragraph and line levels compared to phrases and words (Note that the SCD, defined in Appendix C.5, normalizes the difference of syllable counts by dividing by the number of syllables in the denominator). This indicates that higher syllable counts tend to make matching more challenging. Back-S excels in syllable-related metrics and also achieves the highest BERT-S performance, reflecting its ability to effectively capture the semantics of the input text by directly starting generation after the input text embedding. -values indicate that Back (scratch) excels in syllable-related metrics (SCD, SCErr) and also achieves strong BERT-S performance, reflecting its ability to capture the semantics of the input text effectively.
3.3 Song Form Consistency Evaluation
To verify whether the model differently generates lyrics for different song form types, we measure BERT-S and normalized Levenshtein distance (NLD) between different paragraphs. Note that we do not measure metrics within the same paragraph. Specifically, for each song containing multiple paragraphs, we calculate metrics between distinct paragraphs, excluding comparisons with themselves (e.g., verse1&verse2, verse1&chorus1). This process is performed across all songs, and the average is taken for each case over the total occurrences.
Figure 3 presents the results of generated lyrics by the Back-S model in a square matrix format. As these comparisons are between paragraphs within the same song, we can observe that the BERT-S generally exceeded 0.7. Notably, the BERT-S between the same song form types (diagonal values) are higher than those between different types. For NLD, We can observe results with a similar tendency to those of the BERT-S, with diagonal values being smaller than other values, which denotes the fewer steps to edit the string from one to another. Notably, verses show less similarity among the diagonal values for both metrics. This is expected as verses contain more varied lyrics contents compared to other song form types, but they still exhibit more similarity than the metrics between different song form types. The results suggest that the model tends to produce lyrics that are similar within the same song form and distinct across different forms.
4 Conclusion
Our study presents a sophisticated approach to lyrics generation and infilling, incorporating multi-level granularity for syllable control and effective song form management conditioned on an arbitrary input text. This offers enhanced flexibility in the lyrics generation process. For future work, we plan to extend the capabilities of our models by integrating more flexible control such as genre using genre tags or rhyme control, which we anticipate will further refine the adaptability and creative potential of automated lyrics composition in the music industry.
Limitations
Although our framework is capable of a comprehensive generation of lyrics under various conditions, there are some limitations:
Generation Failures.
There is a small probability of generation failure, where the model may not produce the <END_*> token, resulting in an infinite sequence until it reaches the model’s maximum length.
While generating multiple times can mitigate this issue, it does not guarantee success.
Dependence on Song Form Tagged Datasets.
Our framework requires a dataset with pre-labeled song forms (e.g., verse, chorus).
As a result, the model cannot generate song forms that are not present in the training data, and it cannot be trained on unlabeled, lyrics-only datasets.
Contextual Understanding.
Since our model is not based on large language models (LLMs) and relies solely on SentenceTransformer for capturing meaning, it may struggle to fully grasp the detailed context of the input text, potentially resulting in less nuanced lyric generation.
Ethical Considerations
While our model generates new lyrics based on patterns learned from training data, there is a risk of inadvertently generating content that closely resembles copyrighted lyrics, posing potential copyright infringement issues. Future work should incorporate mechanisms to detect and prevent the generation of lyrics that are too similar to existing copyrighted material.
Additionally, the training data used for our model may contain cultural, gender, or racial biases, which could be reflected in the generated lyrics. Without careful filtering and bias mitigation techniques, the model might inadvertently reinforce harmful stereotypes. It’s important to use diverse, representative, and inclusive datasets, and to explore methods for detecting and correcting bias in lyrics generation.
Appendix A Related Work
Many studies on lyrics generation have attempted to use various features to generate lyrics that match naturally with the melody. Lyrics generation inherits intricate syntactic and semantic challenges from text generation. Notably, within this domain, there has been a significant increase in the utilization of neural networks. Recurrent Neural Networks (RNN) Zaremba et al. (2015), in particular, have garnered acknowledgment for their effectiveness in language modeling and capturing sequential dependencies.
The procedure of converting melodies to expressive lyrics has also been studied, as seen in Chen and Lerch (2020); Ma et al. (2021); Sheng et al. (2021); Lu et al. (2019); Tian et al. (2023); Duan et al. (2024); Zhang et al. (2024). These models adeptly generate lyrics in coherence with the provided melody. For instance, Chen and Lerch (2020) used SeqGANs Yu et al. (2017) for this task, and Ma et al. (2021) proposed AI Lyricists, leveraging MIDI files for lyrics generation. Tailored specifically for melody-to-lyrics generation, models such as Sheng et al. (2021) utilize architectures like MASS Song et al. (2019). Though these models address the problem in various ways, they suffer from a lack of melody-lyrics-aligned dataset. Approaches used in Tian et al. (2023) employ hierarchical frameworks successfully generating lyrics aligned with outlines, trained in an unsupervised manner to address this issue. However, they still do not consider song form or multi-granularity conditioned generation.
Researches focused on syllable count control have employed rule-based and learning-based approaches to produce lyrics that adhere to constraints and convey meaning. Poetry generation is a notable task in this field. For example, studies such as Zhang and Lapata (2014); Yi et al. (2017, 2018) have employed RNN and LSTM for Chinese poetry generation. Additionally, Lu et al. (2019) proposes a method that interprets the alignment between lyrics and melodies as representations of syllable structures. This method uses a multi-channel sequence-to-sequence model that considers both phrasal structures and semantics.
Appendix B Multi-Level Granularity Lyrics Infilling
During lyrics generation, users may need to adjust lyrics to ensure syllabic accuracy or to amend inaccuracies introduced during the initial generation process. To facilitate these corrections while maintaining adherence to the original syllabic constraints, we also propose a syllable-count conditioned lyrics infilling approach, drawing inspiration from Donahue et al. (2020). As in Section 2, this method involves the introduction of special tokens designed to mask lyrics at various levels of granularity (e.g., <INF_P>, <INF_L>, etc.), as demonstrated in our infilling examples shown in Figure 4. We randomly mask parts of the training data by replacing lyrics at different granularities with these tokens, along with their corresponding syllable count tokens. The infilling process begins with the <START> token, signaling the start of infill generation, followed by the injection of song form tokens and the correct answers of the masked lyrics, each accompanied by appropriate <INF_*> and <SYL:> tokens.
At inference, the model follows a similar procedure as Section 2, with the tokens preceding the <START> token serving as the contextual foundation. Special tokens, including song form indicators, infilling directives (e.g. <INF_*>), and syllable counts, are supplied directly to guide the generation, ensuring that each lyrical piece produced matches the specified conditions until the <END_*> is reached.
Appendix C Experiment Setup
C.1 Dataset
For the training and evaluation of our models, we used the Genius Song Lyrics dataset222https://www.kaggle.com/datasets/carlosgdcj/genius-song-lyrics-with-language-information, which comprises approximately 5.1 million multilingual song lyrics, with around 3.3 million lyrics in English. From the dataset, we selectively extracted lyrics featuring explicit song form annotations with each paragraph, aligning with a pre-defined set of song forms: verse, chorus, pre-chorus, post-chorus, and bridge.
C.2 Data Preprocessing
In preprocessing, we implemented several steps to ensure data quality and relevance. This involved the elimination of special characters and the conversion of numerals to words using num2words library333https://pypi.org/project/num2words/. Additionally, we discarded non-lyrics content (e.g., ’guitar solo’ or repetition notations like x2, x3) and lyrics exceeding 500 words. To ensure content suitability and avoid offensive language, we assessed the toxicity of the lyrics using the Detoxify444https://github.com/unitaryai/detoxify library Hanu and Unitary team (2020), excluding any lyrics with a toxicity score above 0.5 to create a more positive and inclusive dataset. The syllable counts for each word in the dataset were calculated using the Syllables555https://pypi.org/project/syllables/ library. The dataset was randomly divided into training, validation, and evaluation subsets. This partition resulted in approximately 340K for training, 18K for validation and 10K for evaluation.
C.3 Data Preparation
Since there is no plan-lyrics paired dataset suitable for our task, we synthesized a generation plan using lyrics samples to simulate the generation plan. In crafting our training samples for the generation task, as illustrated in Figure 2, we employ a strategy of dynamically generating random plans for data sample, guided by a pre-order tree traversal technique, inspired by Donahue et al. (2020). Specifically, we begin by constructing a tree representing the hierarchical structure of the training example, encompassing paragraphs, lines, and words. Through pre-order tree traversal, each subtree is randomly selected with probability of , ensuring a varied selection process. When a subtree is selected, we aggregate the syllable counts across its components and integrate these into the generation plan under the corresponding granularity condition. When a leaf node (word) is selected, there’s a 50% chance we integrate that individual word directly into the generation plan, or a phrase of a random number of words–which ranges from 1 to –is selected for generation plan. Exceptionally, when the selection process is conducted at the line level, all syllables of the selected lines are first combined and displayed as shown in Figure 2. After that, the previously mentioned process is carried out. This approach, as explained in Section 2, serves to inform how many syllables should be placed in a line in advance.
The formation of training samples for the infilling task follows a similar methodology to the generation task, with a key distinction: we mask selected subtrees using a mask token that corresponds to their level of granularity and syllable counts, i.e. <INF_*><SYL:>. The infilling training samples are then formulated by appending the targets of the masked segments alongside with the special tokens, as detailed in Section B and Figure 4.
When the number of tokenized tokens in a sample surpasses the model’s maximum length capacity, we systematically remove paragraphs from the end of the sample until the token count falls below this threshold. This process ensures all training data is compatible with our model’s architectural constraints, facilitating efficient training and optimization.
For the evaluation set, we follow the same input generation plan as used in the training set. Since there are no available input text-lyrics pairs, we created input-lyrics pairs for evaluation by summarizing the lyrics in the evaluation set with a summarization model666https://huggingface.co/facebook/bart-large-cnn Lewis et al. (2019).
@@ C.3 Data Preparation
| / 5.479 | / 75.140 | / 11.744 | / 2.331 | / 0.441 | / 0.856 | / 0.812 | / 0.798 | / 0.850 | / 0.901 | |
C.4 Training and Inference Setup
In all our experiments, we train the GPT-2 model Radford et al. (2019) for 10 epochs, with a batch size of 8 on a single GeForce 3090Ti GPU. We experiment with both training from scratch and using a pre-trained model. For the pre-trained model, we utilize the weights available from HuggingFace777https://huggingface.co/openai-community/gpt2 Wolf et al. (2019). The learning rate is set to 5e-5, with 500 warm-up steps, and training is performed using the AdamW optimizer Loshchilov and Hutter (2017). For all experiments, texts are tokenized using a byte-level Byte Pair Encoding (BPE) with a vocabulary size of 50,257 and an input size of 1024 consecutive tokens, as in the original GPT-2 model. The embedding size of the token embedding was 768 dimensions, matching the embedding size of the SentenceTransformer. Therefore, we directly feed the input text embedding into the GPT-2 model without using any projection layer. That is, we input the continuous sentence embedding vector directly into the model, bypassing the generation model’s embedding layer at the start of the sequence. Therefore, we feed the input text embedding directly into the GPT-2 model without using a projection layer. In other words, the continuous sentence embedding vector is fed directly into the model, bypassing its embedding layer at the start of the sequence.
For inference, our approach employs a sampling strategy, characterized by a top-k value of 20 and a top-p value of 0.9. We calibrated the model’s output diversity and creativity with a temperature of 1.0 and applied a repetition penalty of 1.2. The selection probability in pre-order tree traversal for the construction of training samples is set to 20% for the generation task and 10% for the infilling task.
C.5 Evaluation Metrics
In assessing the performance of our models, we employ a suite of metrics designed to capture various dimensions of model effectiveness, from linguistic predictability to syllabic accuracy and semantic fidelity.
@@ C.5 Evaluation Metrics
Test set perplexity (PPL) measures the ability of model to predict the text within the test set, providing insights into the model’s linguistic prediction capabilities.
We calculated PPL specifically for text tokens, excluding special tokens.
This approach ensures that the focus remains on the model’s ability to predict natural language text, rather than the special tokens that are explicitly provided during the inference process.
In all experimental tables, the PPL values represent the trimmed mean rather than the simple mean, to account for the influence of extreme outliers.
Specifically, the PPL values shown in the tables are the averages calculated after excluding the top 1% of the highest values.
However, the Wilcoxon test was conducted on the complete dataset.
Syllable count distance (SCD) quantifies the discrepancy between the expected set of syllable counts and the syllable counts of the generated text . As defined in Kim et al. (2023b), the calculation is given by:
Syllable count error rate (SCErr) measures how frequently the model generates text with incorrect syllable counts, representing a stricter assessment than SCD by penalizing all inaccuracies equally.
BERT-Score (BERT-S) Zhang et al. (2019) assess semantic coherence between a reference and generated lyrics, by using pre-trained BERT embedding.
In the generation task, BERT-S is computed by comparing the input text with the lyrics produced by the model, whereas in the infilling task, it is calculated by comparing the masked segments with the lyrics that have been infilled in.
Normalized Levenshtein Distance (NLD)
is also adopted to assess the consistency of the song form.
Levenshtein distance, also known as edit distance, measures how many character operations are needed to convert one string into another.
These operations can be deleting, inserting, or replacing.
Since the length of strings varies for each song and song form, we use normalized Levenshtein distance (NLD) Tashima et al. (2018) defined as follows:
, where is the Levenshtein distance and the function calculates the length of the given string.
Our evaluation extends across different levels of granularity, applying each metric to assess both the overall performance (Full) and the performance at specific structural levels within the text. For the ’Full’ evaluation, metrics are computed individually for each sample and then averaged across the dataset. For other granularities, such as paragraph (Para.), it’s possible not all samples will have these specific elements (for instance, a generation plan might not include a generate paragraph step). Therefore, we compute the metrics across the dataset without sample-wise distinction, subsequently averaging these computations to obtain granular insights.

### v2

@@ Abstract
keywords:
lyrics generation, syllable count control1 Introduction
The field of lyrics information processing [1] presents unique challenges that extend beyond traditional text generation, as lyrics must align with both song form structures (such as verses, choruses, and bridges) and specific syllabic constraints. While natural language generation models have shown promise in generating coherent text, applying these models to lyrics introduces additional complexities due to the need for musical and rhythmic alignment.
Several studies have made notable progress in aligning lyrics with melodies through melody-aligned or midi-aligned generation methods [2, 3, 4]. These efforts have been valuable in addressing melody-lyrics synchronization. However, the scarcity of large-scale audio-lyrics aligned datasets still poses a challenge for these approaches, limiting their generalizability across diverse musical forms [5, 6, 7]. Given this data limitation, we concentrate on a crucial aspect of melody-lyrics alignment: the syllable count, which directly impacts the rhythmic fit of lyrics within a musical composition. However, precise syllable control is challenging, as models like GPT [8], which rely on sub-word tokens, may struggle to manage syllables effectively.
@@
2 Related Work
Many studies on lyrics generation have attempted to use various features to generate lyrics that match naturally with the melody. Lyrics generation inherits intricate syntactic and semantic challenges from text generation. Notably, within this domain, there has been a significant increase in the utilization of neural networks.
The procedure of converting melodies to expressive lyrics has also been studied, as seen in [3, 4, 9, 10, 11]. These models adeptly generate lyrics in coherence with the provided melody. For instance, [9] used SeqGANs [12] for this task, and [10] proposed AI Lyricists, leveraging MIDI files for lyrics generation. Tailored specifically for melody-to-lyrics generation, models such as [4] utilize architectures like MASS [13]. Though these models address the problem in various ways, they suffer from a lack of melody-lyrics aligned dataset. Approaches used in [11] employ hierarchical frameworks successfully generating lyrics aligned with outlines, trained in an unsupervised manner to address this issue. However, they do not consider song form or multi-granularity conditioned generation.
Researches focused on syllable count control has utilized rule-based and learning-based approaches to produce lyrics that adhere to constraints and convey meaning. Poetry generation is a notable task in this field. For example, studies such as [14, 15, 16] have employed RNN and LSTM for Chinese poetry generation. Additionally, [3] proposed a method that interprets the alignment between lyrics and melodies as representations of syllable structures. This method uses a multi-channel sequence-to-sequence model that considers both phrasal structures and semantics.
3 Methodology
3.1 Multi-Level Granularity Lyrics Generation
Our system generates lyrics at various granularities (words, phrases, lines, paragraphs) with precise syllable control.
The model is conditioned on input text, song structure (e.g., verse, chorus), and syllable constraints.
Token Structure.
We use a structured token system to align lyrics with song forms, as illustrated in Figure 2(a).
Each section begins with a song form token (e.g., <VERSE>) followed by a syllable count token (<SYL:>), where denotes the designated syllable count for the section, adopting the approach described in [17] and [18].
Inspired by the ideas presented in [19], we implement distinct end tokens for each level of granularity in our model.
For paragraphs, the generation starts with <GEN_P> and ends with <END_P>;
entire lines use <GEN_L> and <END_L>.
For finer segmentation, <GEN_L_NW> tokens are used to indicate the start of such detailed line generation.
Each segment, whether phrases or words, begins with <GEN_N> or <GEN_W> tokens, respectively, and concluded with <END_NW> tokens.
Upon completing a line detailed with phrases and words, an <END_L> token signifies its end.
Semantic Embedding for Input Text.
We employ the SentenceTransformer
[20] model to capture the semantic content of lyrics, a method inspired by the approach of Watanabe et al. [21], who used Stable Diffusion [22] and pre-trained Vision Transformer [23] for semantic extraction in lyrics.
We use semantic embeddings of full lyrics as a substitute condition for the input text during training, enabling the model to generate relevant lyrics by analyzing the semantic content of any input text provided during inference.
Specifically, we incorporate the semantic embedding vector into the initial token embedding of the GPT-2 input.
Inference.
During inference, the model starts with the semantic embedding from the input text and decodes tokens autoregressive.
Special tokens such as song form, syllable counts, and generation directives tokens (<GEN_*>) are provided at decoding step as conditions rather than predicted.
The model continues generating until it produces the appropriate end token (<END_*>), then moves on to the next condition, following the complete generation plan.
3.2 Lyrics Infilling
We also propose a lyrics infilling model that can adjust or correct lyrics based on multi-granularity syllable counts, drawing inspiration from [24]. As in Section 3.1, this method involves the introduction of special tokens designed to mask lyrics at various levels of granularity (e.g., <INF_P>, <INF_L>, etc.), as demonstrated in our infilling examples shown in Figure 2(b). We randomly mask parts of the training data by replacing lyrics at different granularities with these tokens, along with their corresponding syllable count tokens. The infilling process begins with the <START> token, followed by the song form tokens and the correct content for the masked segments, each marked with appropriate infilling and syllable tokens.
During inference, the model follows a similar procedure as Section 3.1, with the tokens preceding the <START> token serving as the contextual foundation and generates the missing parts. Special tokens, including song form indicators, infilling directives (e.g. <INF_*>), and syllable counts, are supplied directly to guide the generation, ensuring that each lyrical piece produced matches the specified conditions until the <END_*> is reached.
@@ 3.2 Lyrics Infilling
Test set perplexity (PPL) measures the ability of model to predict the text within the test set, providing insights into the model’s linguistic prediction capabilities. PPL is computed on text tokens (excluding special tokens) and reported as a trimmed mean (top 1% excluded). Syllable count distance (SCD) [31] quantifies the discrepancy between the expected set of syllable counts and generated counts as:
Syllable count error rate (SCErr) measures how frequently the model generates text with incorrect syllable counts, representing a stricter assessment than SCD by penalizing all inaccuracies equally. BERT-Score (BERT-S) [32] evaluates semantic coherence by comparing the input text with generated lyrics (or masked segments with infilled text) using pre-trained BERT embedding. Normalized Levenshtein Distance (NLD) is also adopted to assess the consistency of the song form. Levenshtein distance, also known as edit distance, measures the number of character operations (deletions, insertions, or substitutions) required to convert one string into another. Normalized Levenshtein distance [33] is defined as:
where is the Levenshtein distance and denotes the string length. Metrics are computed overall (Full) and at specific structural levels (e.g., paragraphs) for detailed insights.